# Peng Cheng, and Jiming Chen

### v2

@@ IV Regret Guarantee
The core of the regret analysis is to construct a uniform policy evaluation bound that holds for all active policies at each stage of the algorithm. In each stage $b$, the active policy set is $\phi_{b}$. Assume that for any $\pi\in\phi_{b}$, the value $V^{\pi}(r,P)$ can be estimated up to an error $\xi_{b}$ by $V^{\pi}(\widetilde{r}^{b},\widetilde{P}^{\text{ref},b})$, then if we eliminate all policies that are at least $2\xi_{b}$ sub-optimal in the sense of $\max_{\pi\in\phi_{b}}V^{\pi}(\widetilde{r}^{b},\widetilde{P}^{\text{ref},b})$, the optimal policy will never be eliminated and all remaining policies will be at most $4\xi_{b}$ sub-optimal. Summing the regret across all stages yields the total regret:
$\text{Regret}(K)\leq 2HL_{1}+\sum_{b=2}^{B}2L_{b}\cdot 4\xi_{b-1}.$ (4)
@@
Uniform bound on $|V^{\pi}(r,P)-V^{\pi}(r,P^{\prime})|$. Next, we aim to bound $\sup_{\pi\in\phi_{b}}\sup_{r^{\prime}}|V^{\pi}(r^{\prime},P)-V^{\pi}(r^{\prime},P^{\prime})|$. This leads to bounding $\sup_{\pi\in\phi_{b}}\mathbb{P}_{\pi}[\mathcal{B}]$, where the event $\mathcal{B}$ occurs when the trajectory visits any infrequently visited tuples in $\mathcal{W}$, which describes the main difference between absorbing MDP and the true MDP. By the definition of $\mathcal{W}$, we can show that these tuples are difficult to visit for any policy in $\phi_{b}$, i.e., with high probability, $\sup_{\pi\in\phi_{b}}\mathbb{P}_{\pi}[\mathcal{B}]\leq\tilde{\mathcal{O}}\left(\frac{X^{3}AH^{4}E_{\varepsilon,\delta}}{L_{b}}\right)$. Using this observation, we bound the bias with proof in Appendix Proof of Lemma 8.
@@
This term bounds how much our refined estimate, $\widetilde{P}^{\text{ref},b}$, deviates from the absorbing MDP, $P^{\prime}$. Using the standard simulation lemma from [31], this variance term can be roughly bounded by $\widetilde{\mathcal{O}}\left(\sqrt{\sum_{h,x,a}\frac{V^{\pi}(1_{h,x,a},P^{\prime})}{N_{h}^{\text{ref},b}(x,a)}}\right)$. Meanwhile, the visitation counts is roughly related to the exploration policy, $N_{h}^{\text{ref},b}(x,a)\propto V^{\pi^{\text{ref},b}}(1_{h,x,a},P^{\prime})$. Thus, the key to bounding this term is our fine exploration policy, $\pi^{\text{ref},b}$. By its design, this policy minimizes the worst-case coverage number $\max_{\pi\in\phi_{b}}\sum_{h,x,a}\frac{V^{\pi}(1_{h,x,a},P^{\prime})}{V^{\pi^{\text{ref},b}}(1_{h,x,a},P^{\prime})}$ over all active policies, with a order of $\mathcal{O}(XAH)$. This ensures sufficient exploration of all relevant state-action pairs, leading to the following guarantee (proof in Appendix Proof of Lemma 9).
@@ V Privacy Guarantee
In this section, we present a secure communication protocol, a shuffle Privatizer adapted from the shuffle binary summation mechanism introduced by [18]. We first describe the mechanism satisfying the Assumption 4, and then show that Algorithm 1, when instantiated with this Privatizer, satisfies the $(\varepsilon,\beta)$-SDP guarantee.
To secure the data of the batch users ($D^{h}$ in Crude Exploration and $D$ in Fine Exploration), we privatize the visitation counts and rewards. Given $n$ users in a batch, we outline the process for a single counter $\widetilde{N}_{h}(x,a,x^{\prime})$ of a specific $(h,x,a,x^{\prime})$. The procedure for other counts like $\widetilde{N}_{h}(x,a)$ and rewards $\widetilde{R}_{h}(x,a)$ is analogous.
Firstly, we allocate privacy budget $\varepsilon^{\prime}=\frac{\varepsilon}{3H}$ and failure probability $\beta^{\prime}=\frac{\beta}{HXA}>0$ to one counter, and define a threshold $\tau=\mathcal{O}(\log(1/\beta^{\prime})/\varepsilon^{\prime 2})$ that controls noise addition. On the local side, each user encodes their data using a randomizer $\mathcal{E}$: if $n\leq\tau$, user $i$ encodes local count as $z_{i}=\mathds{1}_{h}\left(x,a,x^{\prime}\right)+\sum_{j=1}^{m}y_{j}$, where $\{y_{j}\}_{j=1}^{m}$ are i.i.d. sampled, $y_{j}\sim\text{Bernoulli}(1/2)$, and $m=\lceil\frac{\tau}{n}\rceil$; otherwise, the local output is $z_{i}=\mathds{1}_{h}\left(x,a,x^{\prime}\right)+y$ where $y\sim\text{Bernoulli}(\frac{\tau}{2n})$. The user then sends its local message $z_{i}$ to a secure shuffler $\mathcal{F}$, which permutes their messages randomly and then forwards them to the analyzer $\mathcal{G}$.
@@
For the post-processing steps for $\ddot{N}_{h}$, we will adapt the techniques from [15, 29]. Firstly, we solve the linear optimization problem for all $\left(x,a\right)$ below.
@@
The result of our shuffle Privatizer is a set of reliable statistics, with formal guarantees summarized in the following Lemma with proof in Appendix Proof of Lemma 10.
@@
Theorem 11 (Regret Bound under SDP)
@@
VI Numerical Experiment
In this section, we conduct simple numerical experiments to verify our theoretical results for SDP-PE (Algorithm 1).
VI-A Settings
Due to the algorithm’s computational complexity, the simulation is conducted on a simplified “RiverSwim” [32] environment, a standard benchmark with challenging exploration requirements, illustrated in Figure 3. The system consists of four states and two actions “left” and “right”, i.e., $X=4$ and $A=2$. In each episode, the agent starts from the leftmost state and executes its policy. At each step of episode $k$, if the agent chooses the action “left”, it will always succeed. Otherwise, it may fail. It only receives a small amount of reward if it reaches the leftmost state while obtaining a much larger reward if it arrives at the rightmost state. Thus, the MDP naturally requires sufficient exploration to obtain the optimal policy. Each episode is reset every $H=6$ steps.
@@ VI-A Settings
We evaluate the performance of SDP-PE and compare it against the state-of-the-art private algorithm, private UCB-VI under JDP and LDP constraints [29], which updates the policy in every episode, as well as their corresponding non-private versions (PE and UCB-VI). We conduct 20 independent experiments under two privacy regimes $\varepsilon=0.1$ and $\varepsilon=1$, each running $K=2\cdot 10^{4}$ episodes.
@@ VI-B Results and Analysis
Figure 4: Cumulative regret vs. Episode under private RL algorithms
We plot the cumulative regret versus the number of episodes, along with the standard deviation, in Figure 4. Our experimental results offer several key insights that align with our theoretical results.
First, in the non-private setting, UCB-VI, as a theoretically optimal method with per-episode updates, unsurprisingly demonstrates the best performance. Our non-private PE algorithm achieves comparable results, confirming that an effective policy can be learned even with a less frequent, i.e., logarithmic update schedule, and a data-forgetting mechanism.
Second, under both private regimes $(\varepsilon=0.1$ and $\varepsilon=1)$, SDP-PE consistently achieves substantially lower cumulative regret than UCBVI-LDP. This empirically verifies our primary goal of attaining a superior performance-privacy trade-off compared to the local privacy model.
Finally, the experiments also highlight a nuanced trade-off within the regret bounds. While the dominant term of SDP-PE’s regret matches the non-private lower bound and the optimal JDP rate, a gap exists due to the privacy-dependent term’s higher dependency on model parameters $(X,A,H)$. This is visible in the plots, where SDP-PE is slightly higher than that of UCBVI-JDP. This gap reflects the cost of reduced data utilization in our batch-based approach and the absence of a trusted central server.
@@ VII Discussion
Here, we discuss the broader implications of our results in the context of the private learning for the CPS literature.
@@
Performance-privacy trade-off. Our SDP-PE algorithm establishes a favorable regret-privacy trade-off, with a regret bound whose leading term matches the non-private lower bound, $\Omega(\sqrt{XAH^{3}K})$. Crucially, our algorithm significantly outperforms the optimal LDP result [29] by making the privacy costs’ dependency on $1/\varepsilon$ additive –rather than multiplicative– with respect to $K$. The utility matches that of the JDP case, but is achieved within a more resilient, distributed architecture that lacks a single point of failure–a critical feature for robust CPS.
@@
The shuffle model as a resilient privacy architecture. The shuffle model serves as an intermediate trust model between the centralized model (JDP) and the local model (LDP) for privacy-preserving learning in CPS. Firstly, a key theoretical insight is that any RL algorithm satisfying SDP also satisfies JDP via Billboard lemma in [33]. Therefore, the JDP lower bound of $\widetilde{\Omega}(\sqrt{XAH^{3}K}+\frac{XAH}{\varepsilon})$ established by [14], also applies to our setting. It remains an open question whether the lower bound is tight. Secondly, while our protocol involves local noise injection, it does not satisfy formal LDP. Meanwhile, due to the “privacy blanket”, the privacy preservation becomes less costly as the batch size grows and local noise decreases in the later stages.
Computational efficiency for CPS deployment. A key consideration for deploying any control algorithm in a real-world CPS is its computational feasibility. A limitation of SDP-PE is its current computation complexity, as it requires explicitly evaluating a potentially large set of active policies, and solving constrained optimization problems to construct exploration policies in each stage. Future work could improve efficiency by designing, e.g., a constrained policy search module inspired by [28] that implicitly optimize over a constrained policy space rather than maintaining an explicit set.
VIII Conclusion and Future Work

### v3

@@ IV Performance Guarantee
Furthermore, the total number of policy switches during the $K$ episodes is $N_{switch}=\mathcal{O}(H\log K)$.
The bound on the policy switching cost follows directly from the exponential batching schedule and the properties of geometric series. The core of the regret analysis is to construct a uniform policy evaluation bound that holds for all active policies at each stage of the algorithm with proof details in Appendix B-C. In each stage $b$, the active policy set is $\phi_{b}$. Assume that for any $\pi\in\phi_{b}$, the value $V^{\pi}(r,P)$ can be estimated up to an error $\xi_{b}$ by $V^{\pi}(\widetilde{r}^{b},\widetilde{P}^{\text{ref},b})$, then if we eliminate all policies that are at least $2\xi_{b}$ sub-optimal in the sense of $\max_{\pi\in\phi_{b}}V^{\pi}(\widetilde{r}^{b},\widetilde{P}^{\text{ref},b})$, the optimal policy will never be eliminated and all remaining policies will be at most $4\xi_{b}$ sub-optimal. Summing the regret across all stages yields the total regret:
$\text{Regret}(K)\leq 3HL_{1}+\sum_{b=2}^{B}3L_{b}\cdot 4\xi_{b-1}.$ (5)
@@
Uniform bound on $|V^{\pi}(r,P)-V^{\pi}(r,P^{\prime})|$. Next, we aim to bound $\sup_{\pi\in\phi_{b}}\sup_{r^{\prime}}|V^{\pi}(r^{\prime},P)-V^{\pi}(r^{\prime},P^{\prime})|$. This leads to bounding $\sup_{\pi\in\phi_{b}}\mathbb{P}_{\pi}[\mathcal{B}]$, where the event $\mathcal{B}$ occurs when the trajectory visits any infrequently visited tuples in $\mathcal{W}$, which describes the main difference between absorbing MDP and the true MDP. By the definition of $\mathcal{W}$, we can show that these tuples are difficult to visit for any policy in $\phi_{b}$, i.e., with high probability, $\sup_{\pi\in\phi_{b}}\mathbb{P}_{\pi}[\mathcal{B}]\leq\tilde{\mathcal{O}}\left(\frac{X^{3}AH^{4}E_{\varepsilon,\delta}}{L_{b}}\right)$. Using this observation, we bound the bias with proof in Appendix B-A.
@@
This term bounds quantifies the deviation of our refined estimate, $\widetilde{P}^{\text{ref},b}$, deviates from the absorbing MDP, $P^{\prime}$. By employing the standard simulation lemma [dann2017unifying], this variance term can be approximately bounded by $\widetilde{\mathcal{O}}\left(\sqrt{\sum_{h,x,a}\frac{V^{\pi}(1_{h,x,a},P^{\prime})}{N_{h}^{\text{ref},b}(x,a)}}\right)$. These visitation counts are inherently linked to the exploration policy, specifically, $N_{h}^{\text{ref},b}(x,a)\propto V^{\pi^{\text{ref},b}}(1_{h,x,a},P^{\prime})$. Thus, the key to bounding this term lies in the fine exploration policy $\pi^{\text{ref},b}$. By its design, this policy minimizes the worst-case coverage number over all active policies, defined as $\max_{\pi\in\phi_{b}}\sum_{h,x,a}\frac{V^{\pi}(1_{h,x,a},P^{\prime})}{V^{\pi^{\text{ref},b}}(1_{h,x,a},P^{\prime})}$, with a order of $\mathcal{O}(XAH)$. This minimization ensures sufficient exploration of all relevant state-action pairs for policy evaluation, leading to the following guarantee (proof in Appendix B-B).
@@ V Privacy Guarantee
In this section, we present a communication protocol served as a shuffle Privatizer, which is adapted from the shuffle binary summation mechanism introduced by [cheu2019distributed]. We first describe the mechanism satisfying the Assumption 4, and then show that Algorithm 1, when instantiated with this Privatizer, satisfies the $(\varepsilon,\beta)$-SDP guarantee.
For a batch of $n$ users, we outline the procedure for a single counter $\widetilde{N}_{h}(x,a,x^{\prime})$ of a specific $(h,x,a,x^{\prime})$. This procedure applies to batch data $D^{h}$ collected in Crude Exploration or $D$ from Fine Exploration. The processing for other counts like $\widetilde{N}_{h}(x,a)$ and rewards $\widetilde{R}_{h}(x,a)$ is analogous.
Firstly, we allocate privacy budget $\varepsilon^{\prime}=\frac{\varepsilon}{3H}$ and failure probability $\beta^{\prime}=\frac{\beta}{HXA}>0$ to each counter, and define a threshold $\tau=\mathcal{O}(\log(1/\beta^{\prime})/\varepsilon^{\prime 2})$ that controls noise addition. On the local side, each user encodes their data using a randomizer $\mathcal{E}$: if $n\leq\tau$, user $i$ encodes local count as $z_{i}=\mathds{1}_{h}\left(x,a,x^{\prime}\right)+\sum_{j=1}^{m}y_{j}$, where $\{y_{j}\}_{j=1}^{m}$ are i.i.d. sampled from $\text{Bernoulli}(1/2)$, and $m=\lceil\frac{\tau}{n}\rceil$; otherwise, the local output is $z_{i}=\mathds{1}_{h}\left(x,a,x^{\prime}\right)+y$ where yyis sampled from $\text{Bernoulli}(\frac{\tau}{2n})$. Each user then sends its local message $z_{i}$ to a secure shuffler $\mathcal{F}$, which permutes their messages randomly and forwards them to the analyzer $\mathcal{G}$.
@@
For the post-processing steps for $\ddot{N}_{h}$, we will adapt the techniques from [baidifferentially, qiao2023near]. Firstly, we solve the linear optimization problem for all $\left(x,a\right)$ below.
@@
The formal guarantees of our Shuffle Privatizer, leading to a set of reliable statistics, are summarized in the following lemma with proof in Appendix B-D.
@@
Corollary 11 (Regret Bound under SDP)
@@
VI Numerical Experiments
In this section, we present a series of numerical experiments to validate the effectiveness of our proposed SDP-PE framework (Algorithm 1).
VI-A Experimental Setup
The simulation is conducted on a simplified “RiverSwim” [osband2013more] environment, a standard benchmark with challenging exploration requirements, as illustrated in Figure 5. The system consists $X=4$ states, $A=2$ actions (”left”, ”right”), and episode length $H=6$. In each episode, the agent starts at $S_{1}$, receives $r=0.005$ at $S_{1}$, and $r=1.0$ for action ”right” at $S_{4}$. If the agent chooses the action “left”, it will always succeed. Otherwise, it may fail.
@@ VI-A Experimental Setup
We also experiment in a multi-armed bandit (MAB) environment with $A=20$ as a generalization of MDP with $H=1,X=1$. This MAB environment adopts the reward heterogeneity setting considered in [li2024distributed], where the global mean reward is $r^{*}(a)\sim\text{Uniform}(0.0,0.99)$. Each user $k$’s local mean reward is $r_{k}(a)=r^{*}(a)+\xi_{k}(a)$, where $\xi_{k}(a)\sim\mathcal{N}(0,\sigma^{2})$. The user observes a random reward from distribution $\text{Bernoulli}(\text{clip}(r_{k}(a),0,1))$.
We evaluate the performance of SDP-PE and compare it against the state-of-the-art private algorithm, private UCB-VI under JDP and LDP constraints [qiao2023near], which updates the policy in every episode, as well as their corresponding non-private versions (PE and UCB-VI [azar2017minimax]). We conduct 20 independent experiments, each running $K=2\cdot 10^{4}$ episodes. The privacy budget $\epsilon$ is specified in each experiment.
@@ VI-B Results and Analysis
VI-B1 Privacy-Utility Trade-off
Figure 6: Performance evaluation of SDP-PE under RiverSwim environment. (a) Final cumulative regret vs. the privacy budget $\epsilon$. (b) and (c) display cumulative regret vs. time for comparison against baseline algorithms under $\epsilon=0.1$ (high privacy regime) and $\epsilon=1.0$ (low privacy regime), respectively. The shaded area in (b) and (c) indicates the standard deviation.
We first evaluate the cumulative regret under different privacy models in the RiverSwim environment. Figure 6 illustrates the comparison of cumulative regret for SDP-PE against various baselines in the RiverSwim MDP environment with different privacy budgets.
First, the UCBVI-LDP algorithm exhibits the worst regret growth under all privacy regimes. This is because, in RL, LDP requires significant noise to be added at every step, rendering the data almost unusable and making it impractical for real-world network applications. Second, our SDP-PE algorithm achieves substantially lower regret compared to UCBVI-LDP. Its performance closely approaches that of the non-private PE and UCBVI-JDP. This empirically demonstrates that the shuffle model (SDP) provides a superior balance for distributed networks: it avoids the catastrophic utility loss of LDP without requiring a trusted central server (unlike JDP).
Figure 7: Performance evaluation of SDP-PE under MAB environment. ($\varepsilon=1.0,\sigma=0.1$)
In addition, we validate our framework’s generality to bandits. As shown in Figure 7, our SDP-PE algorithm continues to perform excellently in this MAB setting, again surpassing the LDP baseline and approaching the JDP baseline. This confirms our work is a general framework that can be seen as a stateful generalization of the bandit problem, unifying private global optimization for both RL and bandit networks.
Collectively, these results demonstrate that SDP-PE, through the shuffle model of privacy, successfully addresses the privacy bottleneck in online reinforcement learning. It offers an effective solution for distributed network environments, combining robust privacy guarantees with superior learning performance, and showcases its excellent generalization capabilities across different online learning scenarios.
VI-B2 Switching Cost Efficiency
Figure 8: Trade-off between final cumulative regret and policy switch times. ($K=20,000,\varepsilon=1.0$)
Next, we evaluate SDP-PE’s capability in reducing policy deployment cost, a critical engineering challenge for online RL deployments. Figure 8 compares the total number of policy switches across different algorithms over the entire $K=20,000$ episodes. Classic UCBVI algorithms (whether UCBVI, UCBVI-LDP, or UCBVI-JDP), due to their ”per-interaction” policy update mechanism, incur a policy switching cost as high as $\mathcal{O}(K)$, namely $20,000$ total policy updates. In a real-world cellular or cloud network, such frequent ”reconfiguration” operations are untenable and impractical. In contrast, our SDP-PE framework (and its non-private PE variant) leverages exponential batching, reducing the network switching cost to $\mathcal{O}(\log K)$, which is only $88$ policy switches in the experiments. This orders-of-magnitude reduction is attributed to SDP-PE’s ingenious utilization of the shuffle model of privacy’s inherent batching requirement, combined with its policy elimination mechanism. By updating and deploying policies only at the end of each batch, SDP-PE drastically reduces network deployment overhead, making it far more feasible for practical network operations.
These results indicate that SDP-PE is not only superior in terms of privacy but also, by ingeniously leveraging the batching property of SDP, addresses a core engineering challenge of large-scale online RL deployments, significantly enhancing its practical value.
@@ VII Discussion
Here, we discuss the broader implications of our results in the context of the private learning for networked systems literature.
@@
Performance-privacy trade-off. Our SDP-PE algorithm establishes a favorable regret-privacy trade-off, with a regret bound whose leading term matches the non-private lower bound, $\Omega(\sqrt{XAH^{3}K})$ [domingues2021episodic]. Crucially, our algorithm significantly outperforms the optimal LDP result [qiao2023near] by making the privacy costs’ dependency on $1/\varepsilon$ additive, rather than multiplicative with respect to $K$. The utility matches that of the JDP case, but is achieved within a more resilient, distributed architecture that lacks a single point of failure–a critical feature for robust networked systems.
@@
The shuffle model as a resilient privacy architecture. The shuffle model serves as an intermediate trust model between the centralized model (JDP) and the local model (LDP) for privacy-preserving learning. Firstly, a key theoretical insight is that any RL algorithm satisfying SDP also satisfies JDP via Billboard lemma in [hsu2014private]. Therefore, the JDP lower bound of $\widetilde{\Omega}(\sqrt{XAH^{3}K}+\frac{XAH}{\varepsilon})$ established by [vietri2020private], also applies to our setting. It remains an open question whether the lower bound is tight for the SDP setting. Secondly, while our protocol involves local noise injection, it does not satisfy formal LDP. Meanwhile, due to the “privacy blanket” effect, the privacy preservation becomes less costly as the batch size grows and local noise decreases in the later stages.
Policy switching under privacy constraints. The SDP-driven batching strategy in our online RL scenario creates a synergy with the engineering need to reduce “policy switching costs”. In a non-private setting, achieving a near-optimal regret of $\tilde{\mathcal{O}}(\sqrt{K})$ typically requires at least $\Omega(\log\log K)$ switches [qiao2022sample]. However, the necessity of batching under shuffle privacy to overcome aggregated noise leads our SDP-PE algorithm to an $O(\log K)$ switching frequency. This additional switching frequency ($\log K$ vs. $\log\log K$) appears to be necessary to maintain high utility while satisfying strong privacy constraints. Thus, it remains an important open question whether it is possible to achieve fewer policy switches (e.g., at the $\Omega(\log\log K)$ level) while satisfying shuffle privacy and simultaneously maintaining near-optimal regret.
Limitations and future Work. A key limitation of our framework lies in its computational scalability. Our SDP-PE, based on policy elimination, relies on explicitly evaluating a potentially large set of active policies and solving constrained optimization problems to construct exploration policies in each stage. These steps, especially enumerating over all deterministic policies, can lead to exponential time complexity and thus do not scale well to large state-action spaces. A possible direction is to apply softmax (or other differentiable) representation of the policy space and use gradient-based optimization techniques to find approximate solutions. Another promising solution could involve designing a constrained policy search module inspired by [zhang2022near], that implicitly optimize over a constrained policy space rather than maintaining an explicit set.
VIII Related Work
Our work lies at the unique intersection of three research domains: (1) learning in distributed networks, (2) privacy in sequential decision-making, and (3) reinforcement learning.
Learning paradigms in distributed networks. In modern networked systems, learning from distributed entities has become a central challenge. Research in this area can be broadly categorized based on its network structure and interaction paradigm. First, federated learning (FL) [9060868] is a widely studied paradigm characterized by centralized coordination and decentralized data. A central server orchestrates numerous clients, which train models locally and upload updates (rather than raw data) to collaboratively learn a global model. FL typically involves clients uploading data or updates in a periodic, synchronous, batch manner. Secondly, distributed learning [10.1145/3377454], in contrast, lacks a central coordinating node. Instead, individual nodes in the network learn local or shared models through peer-to-peer communication. This paradigm is often more complex, focusing on communication efficiency and the impact of network topology. Thirdly, multi-agent systems (MAS) [10439972], including multi-agent reinforcement learning [ZENG2022119688], focus on problems where multiple autonomous agents interact within a shared environment, primarily studying issues related to game theory among agents. In this paper, we share similarities with FL in its “central learning agent-distributed nodes” structure, yet it differs fundamentally. The distributed nodes act as independent decision-makers and interact with the environment under policies deployed by the central agent, and provide raw (or perturbed) data to it, rather than training models locally in FL or playing games in MAS.
Sequential decision-making in distributed networks. Within the aforementioned networked learning paradigms, sequential decision-making problems, particularly bandits and reinforcement learning, have been extensively studied. Distributed/federated bandits [li2024distributed, 10552083] apply FL principles to bandit problems, aiming to learn a global decision policy from distributed clients’ real-time feedback. For instance, [li2024distributed] solves the problem of learning a global model for global regret minimization in a linear bandit setting, where distributed client feedback interacts in a “parallel batch” manner. Distributed reinforcement learning (DRL) [labbi2025federated] involves distributing RL tasks across multiple computational nodes. However, most DRL work focuses on using multiple parallel agents to accelerate the training of a single task, where agents typically share data or models, without addressing user privacy concerns or unique online task stream characteristics. In contrast to the above synchronous batch models, an “online Streaming Task Sequence” model is also studied within bandits and RL framework [10018417, 8022964]. This model is highly relevant to our setting, where user requests (i.e., episodes) arrive sequentially and dynamically. Typically, only one user interacts with the system at any given time, and the central agent must learn from this real-time, single-user stream.
Online RL algorithms. Regarding online reinforcement learning algorithms, various methods with theoretical guarantees have been extensively investigated. These include algorithms designed for tabular MDPs, such as those based on the Optimism in the Face of Uncertainty principle (e.g., UCBVI [azar2017minimax]), Thompson Sampling (e.g. PSRL [osband2013more]), Policy Elimination (PE) [qiao2022sample] and their variants. For large state spaces, researchers have also developed linear approximation methods [jin2020provably] and policy gradient algorithms [agarwal2021theory] with theoretical guarantees. This paper primarily focuses on tabular MDP environments, utilizing Policy Elimination algorithms as core building blocks, but these must be fundamentally reshaped under shuffle privacy constraint.
Privacy in sequential decision-making. In the “online streaming” model, the feedback uploaded by each user is highly sensitive, making privacy a core challenge. To address this, differential privacy has been widely applied to such sequential decision-making problems. Firstly, JDP (Centralized Privacy) has been studied in bandits [shariff2018differentially] and RL [vietri2020private], JDP requires a fully trusted central server to deal with the raw data, which is an unrealistic assumption in our decentralized, anonymous crowd setting. Meanwhile, LDP (Local Privacy) avoids the trust assumption, but applications in RL [garcelon2021local] and bandits [ren2020multi] show that the immense noise required to protect every trajectory leads to a severe utility collapse (e.g., a multiplicative $\mathcal{O}(\sqrt{K/\epsilon})$ regret term), rendering it impractical for high-utility applications. Recently, SDP (Shuffle Privacy) as an intermediate model, has been successfully applied to bandit problems via action elimination ideas [tenenbaum2021differentially, li2024distributed], proving it can balance trustlessness and high utility. However, to the best of our knowledge, extending the shuffle model to the more complex, online reinforcement learning problem remains an open gap. This work is the first to fill this gap, providing a foundation for shuffle-private episodic RL.
Batched RL. Applying SDP to online RL faces a significant technical challenge, namely, the SDP privacy inherently requires data to be “batched” to achieve privacy amplification. This constraint is non-trivial for the standard online RL model, as it introduces (a) delayed feedback (the agent must wait for a batch) and (b) aggregated noise (the agent observes a noisy aggregate, not a single trajectory), rendering traditional RL algorithms (like UCBVI [azar2017minimax]) ineffective. Concurrently, a separate research area, low-policy-switch RL, also leverages “batching” techniques to solve a purely engineering bottleneck (the $\mathcal{O}(K)$ policy switching cost). These works reduce switching costs to $\mathcal{O}(\log\log K)$ or $\mathcal{O}(\log K)$ [bai2019provably, zhang2022near, johnson2024sampleefficiency]. The core technical innovation of this paper is the discovery and exploitation of the synergy between these two seemingly disparate fields. We demonstrate that the policy elimination framework, used to solve “low-switching-cost”, is the ideal vehicle to solve the “delay and noise” challenges introduced by SDP, especially when modified with our “forgetting” mechanism, and offers a novel paradigm for practical, privacy-preserving online RL in networked systems.
IX Conclusion and Future Work